(5/7) ÷ 6 is equal to: (a) 30/7 (b) 5/42 (c) 30/42 (d) 6/7

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (b) is correct here

= 5/7 ÷ 6/1
= (5/7)/ (6/1)
= (5/7) × (1/6)
= 5/42


Thanks